It says factor f(x) into linear factors given that k is a zero of f(x), f(x)=4x^3+11x^2-115=28 k=4 . Can you help me? My teacher has not gone over this and I do not know how to do it and I have a Final Exam in my class Friday. Thank you very much!!!!!
Answer by rothauserc(4718)

f(x) = 4x^3 +11x^2 -115 = 28, and k = 4 is a zero
:
where is k in f(x)?
:
Let's check x = 4 is a zero of f(x)
:
rewrite f(x)
:
4x^3 +11x^2 -115 -28 = 0
:
4x^3 +11x^2 -143 = 0
:
If we substitute x=4, we should get 0
:
4(4^3) +11(4^2) - 143 = 0
:
256 +176 -143 = 0
:
432 -143 = 0
:
289 does not = 0
:
**********************************
x = 4, is not a zero of f(x)
Do you have f(x) correctly stated?
**********************************
: